# Stability Analysis of Cosmological Perturbations in the Bumblebee Model: Parameter Constraints and Gravitational Waves

### v2

@@ VII Polarization modes of gravitational waves
Figure 1: This figure illustrates the six polarization modes of GWs Eardley:1973zuo . Assuming the propagation direction of GWs is along the “+z+z” direction, these panels show the variation in the relative positions of test particles distributed on a spherical surface. No relative displacement occurs for test particles located along the direction perpendicular to the plane of each figure. The solid and dashed lines represent a phase difference of $\pi$, where the symbol $\rightarrow$” denotes wave propagation to the right, and $\odot$” indicates the wave propagation outward from the page.
@@
$\ddot{h}_{w}-\frac{1}{1-\xi b_{t}^{2}}\partial_{z}\partial_{z}h_{w}=0.$ (141)
According to the relation (139) between the polarization modes and the components of the Riemann tensor, $P_{+}\approx\ddot{h}_{+}$ and $P_{\times}\approx\ddot{h}_{\times}$, the two independent tensor modes ($P_{+}$ and $P_{\times}$) are permitted to propagate. Since tachyonic stability requires $\xi\leq 0$ (77), the speed of the tensor modes is less than the speed of light, with $c_{t}=1/\sqrt{1-\xi b_{t}^{2}}\leq 1$. The equality holds only when the Lorentz-violating parameter $\xi b_{t}^{2}$ is zero.
@@
$\lambda_{p}=\frac{\xi b_{t}}{1-\xi b_{t}^{2}}\zeta_{p},$ (142)
$\ddot{\zeta}_{p}+\left(1+\frac{\xi^{2}b_{t}^{2}}{2\kappa(1-\xi b_{t}^{2})}\right)\delta^{ij}k_{i}k_{j}\zeta_{p}=0.$ (143)
According to the relation (139), $P_{x}\approx-\partial_{z}\dot{\lambda}_{x}$ and $P_{y}\approx-\partial_{z}\dot{\lambda}_{y}$, there are two independent vector modes ($P_{x}$ and $P_{y}$) in the Bumblebee model. Since the Lorentz-violating parameter is very small $|\xi b_{t}^{2}|\ll 1$, the speed of the vector modes exceeds the speed of light. However, if $\xi b_{t}^{2}=0$ (where the propagation speed of the vector perturbations equals the speed of light), GWs of the vector modes are prohibited from propagating, as shown in Eq. (142).
@@
$\phi_{h}=\frac{\kappa+2\xi b_{t}^{2}(\xi-\kappa)}{2\kappa(1-\xi b_{t}^{2})}E,$ (144)
$\ddot{E}+\frac{3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)}{3\kappa(1-\xi b_{t}^{2})}\delta^{ij}k_{i}k_{j}E+\frac{8\kappa}{3\xi^{2}}(1-\xi b_{t}^{2})\overline{V}_{,b^{2}b^{2}}E=0.$ (145)
The relations between the polarization modes and the scalar perturbations are given by $P_{b}\approx\ddot{E}$ and $P_{l}\approx\ddot{E}-2\partial^{2}_{z}\phi_{h}$. Obviously, only one scalar mode is permitted: a mixed mode of $P_{b}$ and $P_{l}$. According to Eq. (109), the propagation speed of the scalar mode exceeds the speed of light. Equation (145) shows that the scalar mode of GWs is massive, $m_{s}=\sqrt{\big(8\kappa(1-\xi b_{t}^{2})\overline{V}_{,b^{2}b^{2}}\big)/(3\xi^{2})}$.
@@
$P_{l}\approx\ddot{E}-2\partial^{2}_{z}\phi_{h}=\ddot{E}+2\frac{\kappa+2\xi b_{t}^{2}(\xi-\kappa)}{2\kappa(1-\xi b_{t}^{2})}\delta^{ij}k_{i}k_{j}E=\mathcal{O}(\xi b_{t}^{2})\delta^{ij}k_{i}k_{j}E=\mathcal{O}(\xi b_{t}^{2})P_{b}.$ (146)
@@
If substituting $\xi b_{t}^{2}=0$ into the scalar perturbation equations (89), one finds that both $E$ and $\phi_{h}$ are constrained to zero. This implies that no scalar mode is permitted under this condition. Especially, if $b_{t}=0$, or $\xi=0$ and $\overline{V}_{,b^{2}b^{2}}=0$, the scalar perturbation equations (89) all reduce to constraints, with only one degree of freedom. Furthermore, if $\xi=0$, $b_{t}\neq 0$ and $\overline{V}_{,b^{2}b^{2}}\neq 0$, there is no scalar perturbation degree of freedom in the Bumblebee model.
@@
Prior to this work, a study on the polarization modes of the Bumblebee theory was conducted in Ref. Liang:2022hxd . The authors considered GWs propagating along the “+z+z” direction with an arbitrary background vector field $b_{\mu}$ in Minkowski background. Their results indicate that the polarization modes of GWs depend on the parameter space, which is divided into three cases: (I) $b_{x}^{2}+b_{y}^{2}\neq 0$ and $b^{t}\neq b^{z}$; (II) $b_{x}^{2}+b_{y}^{2}=0$ and $b^{t}\neq b^{z}$; (III) $b^{t}=b^{z}$. In Case II, which is consistent with our background conditions, five independent modes are presented: $_{+}, P×P_{\times}$, $P_{x}$, and $P_{y}$, along with a massive mixture mode of the $P_{b}$ and $P_{l}$. Our results are consistent with these findings. The distinction lies in the fact that we determine a specific range for the GW speed, which agrees with the results from cosmological perturbations discussed above.
@@
In conclusion, the scale factor $a(t)$ influences the polarizations of GWs, but this effect is very weak compared to that of the perturbations in the cosmological background. In the limit $a\rightarrow 1,\dot{a}\rightarrow 0,\ddot{a}\rightarrow 0$, we find that the Bumblebee model allows the propagation of two tensor modes ($P_{+},P_{\times}$), two vector modes ($P_{x},P_{y}$), and one scalar mode (mixture of $P_{b}$ and $P_{l}$). All GWs of these modes propagate at speeds different from the speed of light. The speed of the tensor modes is less than the speed of light, while the speeds of the vector and scalar modes exceed it. For the mixed scalar mode of $P_{b}$ and $P_{l}$, GWs are massive. But in the small-scale limit or $\overline{V}_{,b^{2}b^{2}}=0$, the mass term for GWs vanishes and the amplitude of $P_{l}$ is significantly smaller than that of $P_{b}$. Especially, if the Lorentz-violating parameter equals zero ($\xi b_{t}^{2}=0$), all vector and scalar modes are not allowed to propagate. In this case, only the tensor modes ($P_{+}$ and $P_{\times}$) propagate at the speed of light. We summarize these results in Table 1. The condition $\xi\leq 0$ is obtained by the tachyonic stability of the vector perturbations.
$\xi b_{t}^{2}\neq 0$
Modes DoF Speeds Mass
Tensor $P_{+},P_{\times}$ 2 $c_{t}<1$ 0
Vector $P_{x},P_{y}$ 2 $c_{v}>1$ 0
Scalar Mixture 1 $c_{s}>1$ $\neq 0$
Table 1: The polarization modes of GWs in the Bumblebee model. Here, we take the limit $a\rightarrow 1,\dot{a}\rightarrow 0,\ddot{a}\rightarrow 0$ and use the condition $\xi\leq 0$.
@@ VIII Conclusion
The formulation of Einstein’s GR marks a qualitative leap in humanity’s understanding of gravitational phenomena. Over the subsequent century, the exploration of gravity has continued unremittingly. In 2015, humanity achieved the first direct detection of GWs LIGOScientific:2016aoc ; LIGOScientific:2016emj , confirming one of the most significant predictions of GR. This milestone ushered in the era of GW astronomy. However, with the deepening of our understanding of the universe, there has been an increasing number of phenomena that GR struggles to explain Smith:1936mlg ; Zwicky:1937zza ; Peebles:2002gy ; tHooft:1974toh ; Goroff:1985th ; Arkani-Hamed:1998jmv ; Randall:1999ee ; Randall:1999vf . Is GR the fundamental theory of gravity? Which theory truly represents the fundamental description of gravity? These questions have become central to contemporary physics.
@@
Through this investigation, we reached the following conclusions. By combining the background field equations with the requirement of cosmic accelerated expansion, the time derivative of the Hubble parameter is constrained to be negative, $\dot{H}<0$. The stability requirement for the vector perturbations indicates that the coupling parameter must be non-positive, $\xi\leq 0$. Combining these two results yields constraints on dark energy and the potential: $w_{D}\leq-1$, $\overline{V}_{,b^{2}}(t_{0})\leq 0$. If $\overline{V}=0$, the background equations indicate that only the cosmological constant $\Lambda$ contributes to the accelerated expansion of the universe, while the Bumblebee field does not. Based on constraint on the speed of tensor modes derived from the GW170817 event and its electromagnetic counterpart GRB170817A, the Lorentz-violating parameter $\xi b^{2}_{t}$ is constrained to the range $6\times 10^{-15}\lesssim\xi b^{2}_{t}\leq 0$. The stability conditions of the scalar perturbations require $\overline{p}_{m,n}>0$. Furthermore, cosmological perturbations and GW analyses indicate that within the Bumblebee model, there exist two independent tensor modes ($P_{+}$ and $P_{\times}$), two independent vector modes ($P_{x}$ and $P_{y}$), and one massive scalar mode (a mixture of the $P_{b}$ and $P_{l}$). The propagation speed of tensor modes is less than the speed of light, while the speeds of vector and scalar modes exceed it. In the small-scale limit or $\overline{V}_{,b^{2}b^{2}}=0$, the amplitude of $P_{l}$ is significantly smaller than that of $P_{b}$, satisfying $P_{l}=\mathcal{O}(\xi b^{2})P_{b}$.
@@
Due to the typically vast distances to GW sources, any deviation in the propagation speed of GWs could result in significant differences in their arrival times. This effect would cause some faint yet significant GW signals to fall outside the windows of data analysis and detection in clear signals Schumacher:2023jxq . Therefore, to test the Bumblebee model, it is crucial to determine the magnitude of its predicted deviations with greater precision and to expand the scope of data analysis. More importantly, detectors with higher sensitivity are necessary to increase the probability of detecting GWs with additional polarization modes. Future space-based GW detectors (Taiji, TianQin, and Lisa projects Luo:2021qji ; TianQin:2015yph ; LISA:2017pwj ) are anticipated to accomplish this task.
@@
Note: At the time of our initial submission, a related work vandeBruck:2025aaa had already been posted on arXiv (arXiv:2509.11647). While both studies explore similar questions, there are several notable differences. The model in Ref. vandeBruck:2025aaa includes an additional coupling term $B_{\mu}B^{\nu}R$ but does not incorporate a cosmological constant $\Lambda$. Additionally, the authors consider a time-dependent background for the Bumblebee field, while we treat it as constant. The analysis in Ref. vandeBruck:2025aaa offers a detailed treatment of the dark energy implications, whereas our study focuses on the connection between perturbations and polarization modes of GWs.
@@ Acknowledgments
We would like to thank Shan-Ping Wu and Chen Tan for useful discussions. This work is supported in part by the National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12475056, No. 123B2074, and No. 12247101), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2025-jdzx07), the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 25JRRA799), the 111 Project (Grant No. B20063), and the Department of Education of Gansu Province: Outstanding Graduate “Innovation Star” Project (Grant No. 2025CXZX-153).

### v3

@@ VII Polarization modes of gravitational waves
Figure 1: This figure illustrates the six polarization modes of GWs Eardley et al. (1973). Assuming the propagation direction of GWs is along the “+z+z” direction, these panels show the variation in the relative positions of test particles distributed on a spherical surface. No relative displacement occurs for test particles located along the direction perpendicular to the plane of each figure. The solid and dashed lines represent a phase difference of $\pi$, where the symbol $\rightarrow$” denotes wave propagation to the right, and $\odot$” indicates the wave propagation outward from the page.
@@
$\ddot{h}_{w}-\frac{1-\sigma b_{t}^{2}}{1-\sigma b_{t}^{2}-\xi b_{t}^{2}}\partial_{z}\partial_{z}h_{w}=0.$ (141)
According to the relation (139) between the polarization modes and the components of the Riemann tensor, $P_{+}\approx\ddot{h}_{+}$ and $P_{\times}\approx\ddot{h}_{\times}$, the two independent tensor modes ($P_{+}$ and $P_{\times}$) are permitted to propagate. Since tachyonic stability requires $\xi\leq 0$ (76), the speed of the tensor modes is less than the speed of light, with $c_{t}=\sqrt{(1-\sigma b_{t}^{2})/(1-\sigma b_{t}^{2}-\xi b_{t}^{2})}\leq 1$. The equality holds only when the Lorentz-violating parameter $\xi b_{t}^{2}$ is zero.
@@
$\lambda_{p}=\frac{\xi b_{t}}{1-(\xi+\sigma)b_{t}^{2}}\zeta_{p},$ (142)
$\ddot{\zeta}_{p}+\left(1+\frac{\xi^{2}b_{t}^{2}}{2\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}\right)\delta^{ij}k_{i}k_{j}\zeta_{p}=0.$ (143)
According to the relation (139), $P_{x}\approx-\partial_{z}\dot{\lambda}_{x}$ and $P_{y}\approx-\partial_{z}\dot{\lambda}_{y}$, there are two independent vector modes ($P_{x}$ and $P_{y}$) in the Bumblebee model. Since the Lorentz-violating parameters are very small $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$, the speed of the vector modes exceeds the speed of light. However, if $\xi b_{t}^{2}=0$ (where the propagation speed of the vector perturbations equals the speed of light), GWs of the vector modes are prohibited from propagating, as shown in Eq. (142).
@@
$\phi_{h}=\frac{\kappa+2\xi b_{t}^{2}(\xi-\kappa)-\kappa\sigma b_{t}^{2}}{2\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}E,$ (144)
$\ddot{E}+\frac{3\kappa+2\xi b_{t}^{2}(\xi-2\kappa)-3\kappa\sigma b_{t}^{2}}{3\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}\delta^{ij}k_{i}k_{j}E+\frac{8\kappa}{3(\xi+2\sigma)^{2}}\big(1-(\xi+\sigma)b_{t}^{2}\big)\overline{V}_{,b^{2}b^{2}}E=0.$ (145)
The relations between the polarization modes and the scalar perturbations are given by $P_{b}\approx\ddot{E}$ and $P_{l}\approx\ddot{E}-2\partial^{2}_{z}\phi_{h}$. Obviously, only one scalar mode is permitted: a mixed mode of $P_{b}$ and $P_{l}$. According to Eq. (120), the propagation speed of the scalar mode exceeds the speed of light. Equation (145) shows that the scalar mode of GWs is massive, $m_{s}=\sqrt{8\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)\overline{V}_{,b^{2}b^{2}}/\big(3(\xi+2\sigma)^{2}\big)}$.
@@
$P_{l}\approx\ddot{E}-2\partial^{2}_{z}\phi_{h}=\ddot{E}+2\frac{\kappa+2\xi^{2}b_{t}^{2}-\kappa b_{t}^{2}(2\xi+\sigma)}{2\kappa\big(1-(\xi+\sigma)b_{t}^{2}\big)}\delta^{ij}k_{i}k_{j}E=\mathcal{O}\big((\xi b_{t}^{2})(\sigma b_{t}^{2})^{0}\big)\delta^{ij}k_{i}k_{j}E=\mathcal{O}\big((\xi b_{t}^{2})(\sigma b_{t}^{2})^{0}\big)P_{b}.$ (146)
@@
From the scalar perturbation equations (88) and the equations of motion (144) and (145), the properties of the scalar GWs depend on the parameter space as follows:
  • Case 1: $\xi b_{t}^{2}\neq 0$. A single massive scalar mode exists, corresponding to the mixed mode of the $P_{b}$ and $P_{l}$. Especially, when $\overline{V}_{,b^{2}b^{2}}=0$, this scalar mode is massless and superluminal.
  • Case 2: $\xi=0,\sigma b_{t}^{2}\neq 0$. A single massive scalar mode is allowed, corresponding to the mixed mode of the $P_{b}$ and $P_{l}$. Especially, when $\overline{V}_{,b^{2}b^{2}}=0$, this reduces to a massless $P_{b}$ mode propagating luminally.
  • Case 3: $\xi b_{t}^{2}=\sigma b_{t}^{2}=0$. No scalar GWs exist in this regime. In particular, when $\xi=\sigma=0,b_{t}\neq 0,\overline{V}_{,b^{2}b^{2}}\neq 0$, there is no scalar perturbation degree of freedom in the Bumblebee model. In all other situations satisfying the case condition, a single scalar degree of freedom remains, originating from the Bumblebee field.
@@
Prior to this work, a study on the polarization modes of the Bumblebee theory was conducted in Ref. Liang et al. (2022) where $\sigma=0$. The authors considered GWs propagating along the “+z+z” direction with an arbitrary background vector field $b_{\mu}$ in Minkowski background. Their results indicate that the polarization modes of GWs depend on the parameter space, which is divided into three cases: (I) $b_{x}^{2}+b_{y}^{2}\neq 0$ and $b^{t}\neq b^{z}$; (II) $b_{x}^{2}+b_{y}^{2}=0$ and $b^{t}\neq b^{z}$; (III) $b^{t}=b^{z}$. In Case II, which is consistent with our background conditions, five independent modes are presented: $_{+}, P×P_{\times}$, $P_{x}$, and $P_{y}$, along with a massive mixture mode of the $P_{b}$ and $P_{l}$. Our results are consistent with these findings. The distinction lies in the fact that we determine a specific range for the GW speed, which agrees with the results from cosmological perturbations discussed above.
@@
In conclusion, the scale factor $a(t)$ influences the polarizations of GWs, but this effect is very weak compared to that of the perturbations in the cosmological background. In the limit $a\rightarrow 1,\dot{a}\rightarrow 0,\ddot{a}\rightarrow 0$, we find that the Bumblebee model allows the propagation of two tensor modes ($P_{+},P_{\times}$), two vector modes ($P_{x},P_{y}$), and one scalar mode (the $P_{b}$ or the mixture of the $P_{b}$ and $P_{l}$). All GWs of these modes propagate at speeds different from the speed of light when $\xi b_{t}^{2}\neq 0$. The speed of the tensor modes is less than the speed of light, while the speeds of the vector and scalar modes exceed it. For the mixed scalar mode of $P_{b}$ and $P_{l}$, GWs are massive. However, under either the small-scale limit or $\overline{V}_{,b^{2}b^{2}}=0$, the mass term for GWs vanishes. In these regimes, the amplitude of $P_{l}$ becomes significantly smaller than that of $P_{b}$. Furthermore, $P_{l}$ vanishes entirely if $\xi=0,\sigma b_{t}^{2}\neq 0$. In particular, when the Lorentz-violating parameters are zero ($\xi b_{t}^{2}=\sigma b_{t}^{2}=0$), all vector and scalar modes are non-propagating. In this case, only the tensor modes ($P_{+}$ and $P_{\times}$) propagate at the speed of light. We summarize these results in Table 1. The condition $\xi\leq 0$ is obtained by the tachyonic stability of the vector perturbations.
Case Conditions Type Modes d.o.f. Speed Mass
Tensor $P_{+},P_{\times}$ 2 <1<1 0
Case 1 $\xi b_{t}^{2}\neq 0$ Vector $P_{x},P_{y}$ 2 >1>1 0
Scalar Mixture 1 >1>1 $\neq 0$
Tensor $P_{+},P_{\times}$ 2 11 0
Case 2 $\xi=0,\sigma b_{t}^{2}\neq 0$ Vector no 0 - -
[…]
Tensor $P_{+},P_{\times}$ 2 11 0
Case 3 $\xi b_{t}^{2}=\sigma b_{t}^{2}=0$ Vector no 0 - -
Scalar no 0 - -
Table 1: The polarization modes of GWs in the Bumblebee model. Here, we take the limit $a\rightarrow 1,\dot{a}\rightarrow 0,\ddot{a}\rightarrow 0$ and use the condition $\xi\leq 0$. “d.o.f.” represents the degrees of freedom in the propagation of GWs. The speeds of the scalar modes represent the speed of the scalar GWs when the mass term is ignored. In case 2, the $P_{l}$ component within the mixed mode of the $P_{b}$ and $P_{l}$ vanishes if $\overline{V}_{,b^{2}b^{2}}=0$.
@@ VIII Conclusion
The formulation of Einstein’s GR marks a qualitative leap in humanity’s understanding of gravitational phenomena. Over the subsequent century, the exploration of gravity has continued unremittingly. In 2015, humanity achieved the first direct detection of GWs Abbott and others (2016b, a), confirming one of the most significant predictions of GR. This milestone ushered in the era of GW astronomy. However, with the deepening of our understanding of the universe, there has been an increasing number of phenomena that GR struggles to explain Smith (1936); Zwicky (1937); Peebles and Ratra (2003); ’t Hooft and Veltman (1974); Goroff and Sagnotti (1986); Arkani-Hamed et al. (1998); Randall and Sundrum (1999a, b). Is GR the fundamental theory of gravity? Which theory truly represents the fundamental description of gravity? These questions have become central to contemporary physics.
@@
Through this investigation, we reached the following conclusions. By combining the background field equations with the requirement of cosmic accelerated expansion, the time derivative of the Hubble parameter is constrained to be negative, $\dot{H}<0$ (29), and a constant $b_{t}$ implies $\sigma\neq-\tfrac{1}{2}\xi$ (27). The stability requirement for the vector perturbations indicates that the coupling parameter must be non-positive, $\xi\leq 0$ (76). Combining these two results yields constraints on dark energy: $w_{D}\leq-1$ if $\sigma$ vanishes (36). The stability conditions of the scalar perturbations show that the absence of ghost instability can be frequency-dependent (104). To ensure the Bumblebee model is ghost-free at all perturbation frequencies, the ghost-free condition must reduce to the form given in Eq. (105) (which, under the assumption $|\xi b_{t}^{2}|,|\sigma b_{t}^{2}|\ll 1$, requires $b_{t}^{2}\lesssim-\tfrac{2H^{2}}{\kappa\dot{H}}$). Based on the analysis combining the tensor, vector, and scalar perturbations, we found that there are seven perturbative degrees of freedom in the cosmological background, comprising two tensor, two vector, and two scalar modes, plus one in the matter sector. To analyze the stability and propagation characteristics of GWs, we considered the small-scale limit. Based on the constraint on the speed of tensor modes derived from the GW170817 event and its electromagnetic counterpart GRB170817A, the Lorentz-violating parameter $\xi b^{2}_{t}$ is constrained to the range $6\times 10^{-15}\lesssim\xi b^{2}_{t}\lesssim 0$ (48). For the scalar modes in the small-scale limit, the Laplacian stability requires $\overline{p}_{m,n}>0$ (117). Furthermore, we investigated the polarization modes of GWs in the Minkowski limits. The results indicate that within the Bumblebee model, there exist two independent tensor modes ($P_{+}$ and $P_{\times}$), two independent vector modes ($P_{x}$ and $P_{y}$), and one scalar mode (the $P_{b}$ or the mixture of the $P_{b}$ and $P_{l}$), see Table 1. The propagation speed of tensor modes is less than the speed of light when $\xi b_{t}^{2}\neq 0$ (141), while the speeds of vector (143) and scalar (145) modes exceed it.
@@
Due to the typically vast distances to GW sources, any deviation in the propagation speed of GWs could result in significant differences in their arrival times. This effect would cause some faint yet significant GW signals to fall outside the windows of data analysis and detection in clear signals Schumacher et al. (2023). Therefore, to test the Bumblebee model, it is crucial to determine the magnitude of its predicted deviations with greater precision and to expand the scope of data analysis. More importantly, detectors with higher sensitivity are necessary to increase the probability of detecting GWs with additional polarization modes. Future space-based GW detectors (Taiji, TianQin, and Lisa projects Luo et al. (2021); Luo and others (2016); Amaro-Seoane and others (2017)) are anticipated to accomplish this task.
@@
Note: At the time of our initial submission, a related work van de Bruck et al. (2025) had already been posted on arXiv (arXiv:2509.11647). While both studies explore similar questions, there are several notable differences. The model in Ref. van de Bruck et al. (2025) includes an additional coupling term $B_{\mu}B^{\nu}R$ (Following the reviewer’s suggestion, the coupling term has been added to our latest revision.) but does not consider a separate cosmological constant term $\Lambda$. Instead, the potential term $V$ plays the role of a cosmological constant. Additionally, the authors consider a time-dependent background for the Bumblebee field, while we treat it as constant. The analysis in Ref. van de Bruck et al. (2025) offers a detailed treatment of the dark energy implications, whereas our study focuses on the connection between perturbations and polarization modes of GWs.
@@ Acknowledgments
We would like to thank Shan-Ping Wu and Chen Tan for useful discussions. This work is supported in part by the National Key Research and Development Program of China (Grant No. 2021YFC2203003), the National Natural Science Foundation of China (Grants No. 12475056, No. 123B2074, and No. 12247101), Gansu Province’s Top Leading Talent Support Plan, the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2025-jdzx07), the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 25JRRA799), the 111 Project (Grant No. B20063), and the Department of Education of Gansu Province: Outstanding Graduate “Innovation Star” Project (Grant No. 2025CXZX-153).